OpenStudy (anonymous):

In a bag of marbles, you have three green, two yellow, six blue, and nine red marbles. What is the probability of picking a yellow marble from the bag?

OpenStudy (anonymous):

\[ \frac{\#\ yellow}{\#\ total\ marbles} \]

OpenStudy (anonymous):

2/20 = 1/10 there is a 10% or 1/0 chance.

OpenStudy (anonymous):

1/10